Fermi National Laboratory


Accelerator Update

Friday November 7 through Monday November 10
All shifts began in shutdown.

On Friday, cryo system technicians began cooling the Tevatron (TeV) sector F1, B1, and D1. The P2 transfer line lost its vacuum to atmosphere; experts investigated and later repaired the problem. Late on the day shift, Operations received permission to run beam in the Main Injector (MI). Communication problems held off Linac beam for about an hour.

On Saturday, the midnight shift discovered the MiniBooNE Horn tripped off. It would reset, but would not stay on. Operators left the Horn off until experts could investigate. Around 4:30 AM, an operator and a MI expert discovered and repaired a MI ground fault. By 7 AM Operations had beam circulating in MI. Experts repaired the MiniBooNE Horn around mid afternoon. By 4 PM, Operations had accelerated MI beam to 120 GeV. Later, MI experts conducted damper studies. At 10:30 PM feeder 45 tripped, knocking off the TeV and causing a power glitch that hit Linac, Booster, and CHL. Operations called in many experts to help with recovery.

On Sunday, operators discovered a major LCW leak in a copper line in the TeV sector F1. The leak was valved out and will be repaired on Monday. The Duty Electrician had the power restored to the TeV by 3 AM. The fault is in between A1 and A2. At 5:30 AM, the Cryo Coordinator reported that all cryo compressors had been recovered and that all cryo systems were in cooldown mode expect for A1, A2, and A3. He said that they lost 1/3 to 1/2 of the helium inventory. Operations re-established beam in Linac at 7 AM and beam to Booster by 7:30 AM. At 10 AM operators found another LCW leak at TeV sector B4. For those curious minds, here are pictures of what the operators found. Water techs repaired these leaks later in the afternoon.

MI experts accelerated beam to 150 GeV at 3 PM.

Future Plans:
Techs will pull a new high voltage cable between TeV sectors A1 and A2. This means that power will remain off at TeV sectors A1, A2, and A3 through to Wednesday. Cryo will receive a helium supply on Wednesday. These houses and the rest of the TeV should be cold by Thursday. The schedule has been jumbled by this feeder problem, but the accelerator complex is still set to begin startup this coming weekend.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/10/2003   email Fermilab